1. Chuck has the following quasi-linear utility function: a) Derive Chucks demand curve for x as a function of P,and P b) Derive Chucks demand for for y c) Is y a normal good?
Answer #1

according to the question:

a)

subject to where pr and pyare the prices of z and y respectively

2 =1-Ny = 0 (2)

i)L

Using equations (1) and (2), we get

rac{1}{2 sqrt x} = lambda p_x rac{1}{p_y} = lambda rac{1}{2 sqrt x} = rac{1}{p_y} p_x x = (rac{p_y}{2 p_x})^2

Demand function of x :

וצ2

It depends on the prices of y and x. Since its a quasilinear function in x so demand of x does not depend on income.

b) Demand function of y:

Using the budget constraint equation and demand function of x :

and (Py) Apr Py 4pr

Thus, demand for y depends on prices as well as on income.

c) A normal good is a good whose demand increases as the consumer's income increases.

To check whether y is a normal good or not, we have to evaluate the income elasticity of y.

M Py dy

Income elasticity = % change in demand / % change in M

Edy /

1-p E,

Em is positive since M, y and p_y are positive numbers. So, y is the normal good.